OpenStudy (superhelp101):

If 6.5 L of water is added to 4.5 L of a 2.8 M Ca(OH)2 solution, what is the molarity of the new solution? 0.25 M 0.62 M 1.1 M 1.9 M

OpenStudy (superhelp101):

@Somy

OpenStudy (somy):

write a balanced reaction

OpenStudy (somy):

wait sorry it looks like dilution

OpenStudy (somy):

\( C_1 \times V_1\) = \( C_2 \times V_2\)

OpenStudy (superhelp101):

what is the C and V stand for?

OpenStudy (somy):

C= molarity V= volume

OpenStudy (superhelp101):

i got 1.1M?

OpenStudy (somy):

how?

OpenStudy (somy):

\(C_1 \times 6.5L\) = \(2.8M \times 4.5L\)

OpenStudy (superhelp101):

1. M = 12.6/(6.5+4.5) 2. M = 12.6/11 3. M = 1.145

OpenStudy (asib1214):

you should rearrange the formula so you know what your are solving for.... it should look like this....